OpenStudy (anonymous):

How do I find the limx->0 of ln((1+x)^(5/x)). I think it has something to do with L'Hop's rule but I can't see how it is 0/0 of infinity/infinity. Thanks :)

OpenStudy (freckles):

Bring the 5/x down

OpenStudy (freckles):

\[\lim_{x \rightarrow 0} \frac{5 \ln(1+x)}{x}\]

OpenStudy (freckles):

you have 0/0

OpenStudy (anonymous):

for some reason the equation is just coming up as characters on my computer...can you write it out using the draw tool for me please :)

OpenStudy (anonymous):

refresh your browser

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

any idea what \[\lim_{x\to 0}(1+x)^{\frac{5}{x}}\] is?

OpenStudy (anonymous):

well I would have made it \[x \sqrt{(1+x)^{5}} \] and so them the lim x-> 0 would be 0. Right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh...um...help?! :)

OpenStudy (anonymous):

take urp @freckles suggestion

OpenStudy (anonymous):

But I don't understand how they bought it down. What rule is that?

OpenStudy (freckles):

power rule for log

OpenStudy (freckles):

\[\ln(x^r)=r \ln(x)\]

OpenStudy (anonymous):

oh...oops...forgot about that one :) so if we did that... \[\frac{ 5*\ln(1+0) }{ 0 }\] \[\frac{ 0 }{ 0}\]

OpenStudy (anonymous):

because ln(1) = 0 yes?

OpenStudy (freckles):

yes ln(1) is 0